# Topological phase transition through tunable nearest-neighbor interactions in a one-dimensional lattice

### v1

@@ Abstract
We investigate the phase diagram of a one-dimensional model of hardcore bosons or spinless fermions with tunable nearest-neighbor interactions. By introducing alternating repulsive and attractive interactions on consecutive bonds, we show that the system undergoes a transition from a bond-ordered (BO) phase to a charge-density wave-II (CDW-II) phase as the attractive interaction strength increases at a fixed repulsive interaction. For a specific interaction pattern, the BO phase exhibits topological properties, which vanish when the pattern is altered, leading to a transition from a topological BO phase to a trivial BO phase through a gap-closing point where both interactions vanish. We identify these phases using a combination of order parameters, topological invariants, edge-state analysis and Thouless charge pumping. By extending our analysis beyond half-filling, we explore the phase diagram across all densities and identify the superfluid (SF) and the pair-superfluid (PSF) phases, characterized by single-particle and bound-pair excitations at incommensurate densities. The proposed model is experimentally realizable in platforms such as Rydberg excited or ultracold atoms in optical lattices, offering a versatile framework to study such interplay between topology and interactions in low-dimensional systems.
@@ I Introduction
The study of topological phases and phase transitions has been a subject of intense research in condensed matter physics Hasan and Kane (2010). Primarily based on a single-particle picture, the topological phases or topological insulators have shaped our understanding of solid-state physics by introducing a new framework for classifying condensed matter systems. The central feature that distinguishes topological phases from non-topological ones is the bulk-edge correspondence, which links the bulk topological invariants to the existence of robust edge or interface states at the boundaries between topologically distinct phases. These edge states are protected by certain underlying symmetries and remain stable against small perturbations until the bulk gap remains finite. Such phases are known as the symmetry-protected topological (SPT) phases Senthil (2015a); Rachel (2018). The robustness of the SPT phases makes them highly promising for applications in fields like topological quantum computing, scattering-free wave transport, and several other advanced technologies von Klitzing (2017); Senthil (2015b); Fidkowski and Kitaev (2011); Pollmann et al. (2010); Gu and Wen (2009); Chen et al. (2012); Pachos and Simon (2014).
@@
One of the celebrated models that possesses such a non-trivial gapped SPT phase is the one-dimensional Su-Schrieffer-Heeger (SSH) model Su et al. (1979) where electrons experience dimerized nearest-neighbor hoppings. The appearance of such unconventional phenomenon in a simple one-dimensional setting has led to numerous exploration of the topological phases in systems described by the SSH model and its variants Asbóth et al. (2016a); Anastasiadis et al. (2022); Li et al. (2014); Pérez-González et al. (2019); Ahmadi et al. (2020). Recent progress in the experimental front has propelled the observation of the topological phases in various quantum simulators such as cold atoms in optical lattices, photonic lattices, superconducting circuits, trapped ion arrays, mechanical systems, acoustic systems, electrical circuits, and Rydberg atomic set-ups Atala et al. (2013); Nakajima et al. (2016); Lohse et al. (2016); Mukherjee et al. (2017); Lu et al. (2014); Thatcher et al. (2022); Meier et al. (2016); Xie et al. (2019); Kitagawa et al. (2012); Leder et al. (2016); de Léséleuc et al. (2019); Le et al. (2020); Jörg et al. (2024).
@@
Topological phases in non-interacting systems have yielded a wealth of insights and established a robust framework based on band topology and topological invariants. However, interacting systems provide a more challenging domain owing to the competing effects of symmetries, strong correlations, lattice topology, and particle statistics Rachel (2018). This has motivated numerous studies which have explored the effects of inter-particle interactions on the stability of topological phases arising from the SSH-type models Li et al. (2013); Taddia et al. (2017); Grusdt et al. (2013); Manmana et al. (2012); Yoshida et al. (2018); Ryu and Hatsugai (2002); Di Liberto et al. (2016, 2017); Nakagawa et al. (2018); Fraxanet et al. (2022); Wang et al. (2015); Barbiero et al. (2018); Ye et al. (2016); Kuno (2019); Montorsi et al. (2022); Julià-Farré et al. (2024); Ryu and Hatsugai (2002); Delplace et al. (2011); Zhou et al. (2023). Interestingly, recent studies have shown that strong interactions under proper conditions can favor topological phases  Li and Haldane (2008); Julià-Farré et al. (2022); Argüello-Luengo et al. (2024a); Dalla Torre et al. (2006); Rossini and Fazio (2012); Berg et al. (2008); Greschner et al. (2020); Padhan et al. (2024). Even more fascinating is a scenario where interactions alone induce topological phases in systems that are completely trivial in the non-interacting limit Mondal et al. (2022a); Parida et al. (2024); Nigam et al. (2024). Recently, several such interaction effects and the interaction-induced SPT phases have been observed using various experimental platforms Sompet et al. (2022); Li et al. (2023); de Léséleuc et al. (2019); Jürgensen et al. (2023, 2021).
@@
Among the existing experimental platforms the systems of ultracold dipolar quantum gases in optical lattices have emerged as versatile platforms for controlling and manipulating the requisite off-site inter-particle interactions which appropriately mimic the models that host topological phases. By exploiting the tunability of such interactions, we reveal in this work that in one dimensional lattice, an interaction-induced SPT phase can emerge at half-filling of hardcore bosons or spinless fermions. By assuming uniform hopping amplitudes of the particles (i.e., no dimerization of the hopping) and allowing attractive and repulsive NN interactions on the alternate bonds of the lattice as depicted in the schematic in Fig. 1(a), we show that the system turns topological for any finite values of the interaction strengths, while the non-interacting limit is a gapless superfluid (SF) or Luttinger liquid (LL). This allows for a phase transition from topological to trivial phase by altering the pattern of the NN interaction (i.e., by changing from attractive-repulsive to repulsive-attractive as depicted in the schematic in Fig. 1(b)) through the point when both the interactions vanish. We demonstrate the signatures of the topological phase and the phase transition through various equilibrium properties. Additionally, we identify further evidence of the topological phase transition via Thouless charge pumping. For completeness, we extend our analysis to explore the situation at all the densities. Finally, we discuss the experimental feasibility of the model.
@@
Figure 1: Schematic illustration of the interaction pattern in NN bonds of the lattice, where $+$ denotes repulsive interaction and -- denotes attractive interaction.
@@
The structure of this paper is as follows. In Sec. II, we describe the model under consideration and the methodology used to obtain all the results. Sec. III discusses the key findings, focusing on two different natures of the interaction strengths in two nearest-neighbor bonds. Finally, Sec. IV summarizes the main results. The details of the analytical calculations are presented in the Appendix.
@@ II Model and Method
We consider a Hamiltonian for a system of interacting hardcore bosons with dimerized NN interactions as given by
[…]
@@
where $i$ denotes the site index, $a_{i}$ ($a^{\dagger}_{i}$) are the annihilation (creation) operators at site $i$, and $n_{i}=a_{i}^{\dagger}a_{i}$ represents the number operator at site $i$. The parameter $t$ corresponds to the hopping amplitude between NN sites, while $V_{1}$ and $V_{2}$ denote the strengths of the alternating NN interactions. In our numerical calculations presented below, we set $t=1$ which sets the energy scale in the system.
@@
[…]
@@
The Hamiltonian in Eq. (3) can also be mapped to a model of interacting spinless fermions using Jordan-Wigner transformation Jordan and Wigner (1928) defined as
@@
We have assumed open boundary conditions for this mapping, as will be used throughout this work. With periodic boundary conditions, however, an additional boundary term arises, which depends on the total number of fermions (or total $S_{z}$) in the chain. The above mappings ensure that the models shown in Eqs. 1, 3 and 5 share similar ground state phase diagrams. The results presented below are all obtained by considering the model shown in Eq. 1 which corresponds to a system of hardcore bosons.
The above model has been previously studied using mean-field theory Fisher et al. (1989); Sheshadri et al. (1993); Jaksch et al. (1998); van Oosten et al. (2001), the density-matrix renormalization group (DMRG) White (1992); Schollwöck (2005) method as well as the bosonization approach, for both hardcore bosons and spins  Mondal et al. (2022a); Nigam et al. (2024). These findings highlight that while the non-interacting version of the model remains gapless, introducing a dimerized NN repulsion for hardcore bosons (or dimerized repulsive ZZ exchange interaction for spins) leads to a topological phase and an associated topological phase transition. The topological phase transition occurs as a function of the interaction dimerization which is tuned by fixing the NN interaction strengths for the even (odd) bonds and varying it on the odd (even) bonds of the lattice. This phase transition is marked by a gap-closing point, occurring when the interactions are of equal strengths which guarantees translational symmetry on the lattice. However, the scenario considered in this work is completely different as we assume the NN interaction to be attractive in nature in either all the even or odd bonds (see Fig. 1).
@@ III Results
In this section, we discuss our findings in detail. We first present the bulk phase diagram and then analyze the topological properties at half-filling. We also provide the signature of the topological phase through Thouless charge pumping. Subsequently, we analyze the scenario by moving away from half-filling. In the end, we briefly discuss the experimental feasibility of the model.
Figure 2: (a) The phase diagram of the model in the $V_{1}-V_{2}$ plane at half-filling shows the BO0, BOπ and the CDW-II phases. The red circle at the origin depicts the gapless critical point. The black solid diamonds indicate the boundary separating the BO and the CDW-II phases. (b) shows a cartoon diagram depicting the interaction pattern considered. (c), (d) and (e) show the schematics for the BO0, CDW-II and BOπ phases respectively.
@@ III.1 Phase diagram at half-filling
The ground state phase diagram of the system of hardcore bosons at half-filling is depicted in Fig. 2(a) in the $V_{1}$-$V_{2}$ plane. The upper left (lower right) corner of the figure corresponds to the situation when $V_{1}<0$, $V_{2}>0$ ($V_{1}>0$, $V_{2}<0$). For each case, we numerically find that the entire phase diagram is gapped and there is a gapped-to-gapped transition. In the following, we identify the two gapped phases as the bond order (BO) and the charge-density wave (CDW) phases except at the origin where the system is gapless (denoted by a red circle). We find that for any finite value of $V_{2}$ when $V_{1}=0$, the ground state is a BO phase (the upper left part of the phase diagram). However, as $V_{1}$ is increased on the attractive side, a transition to a CDW phase occurs. This results in a gapped-to-gapped transition as a function of attractive $V_{1}$ for each value of repulsive $V_{2}$. A similar situation is found upon interchanging the nature of the interactions i.e., by making $V_{1}$ repulsive and $V_{2}$ attractive; the corresponding phase diagram is shown in the bottom right corner of Fig. 2(a). In the following, we quantify these phases by computing different relevant physical quantities as a function of $V_{1}=-V_{2}=V$ chosen for simplicity. This corresponds to an exemplary cut drawn in the phase diagram which is denoted by a dashed line in Fig. 2(a). Along this line, when $V<0$, all the odd bonds experience attractive interaction while the even bonds experience repulsive interaction due to the boundary condition (OBC). Conversely, for $V>0$, the odd bonds become repulsive, and even bonds become attractive.
Figure 3: Single-particle excitation gaps $G_{1}$ and $G_{2}$ plotted as a function of $V$ along the dashed line drawn in Fig. 2. $G_{2}$ (solid curve) remains finite in both the gapped phases. $G_{1}$ remains finite in the BO0 and CDW-II phases and vanishes in the BOπ phase due to the existence of zero-energy edge states. The insets shows the finite-size extrapolation of gaps $G_{1}$ (lower right corner) and $G_{2}$ (upper left corner) in three different parameter regimes corresponding to the gapless phase ($V=0.0$), BO phase ($V=-2.0$) and the CDW phase ($V=-6.0$). In the inset, the solid lines are the fitted curve and the symbols represent the numerical data.
Figure 4: The correlation function $\Gamma(r)$ plotted as a function of distance ($r=|i-j|$) showing power-law and exponential decays for $V=0.0$ and $V=-2.0$ respectively.
@@
$\mu^{+}=E_{N+1}-E_{N}~{}~{}~{}\text{and}~{}~{}~{}\mu^{-}=E_{N}-E_{N-1},$ (7)
are the chemical potentials corresponding to adding and removing a particle respectively in the system, and $E_{N}$ represents the ground state energy of the system with $N$ bosons. In Fig. 3, we plot the extrapolated values of the gap $G_{1}$ (dotted plus curve) as a function of $V_{1}=-V_{2}=V$. It can be seen that the gap remains finite in the region $V>0$. However, in the region $V\leq 0$, the gap $G_{1}$ vanishes up to a critical $V\sim-4.3$, after which it becomes finite. In the following, we will show that while the vanishing of $G_{1}$ at $V=0$ is due to the gapless nature of the system at that point, for $-4.3\lesssim V<0$, the zero gap is due to the topological nature of the system. Interestingly, when the gap to the next single-particle excitation ($G_{2}$) is calculated using the formula
$G_{2}=\mu^{+}_{n}-\mu^{-}_{n},$ (8)
where
$\mu^{+}_{n}=E_{N+2}-E_{N+1}~{}~{}~{}\text{and}~{}~{}~{}\mu^{-}_{n}=E_{N-1}-E_{N-2},$ (9)
we find that the system is gapped for all the values of $V\neq 0$ (red solid curve). The insets in Fig. 3 depict the finite-size extrapolation of $G_{1}$ and $G_{2}$ for different values of $V$ (see the figure caption for details). At the gapless point $V=0$, the system exhibits off-diagonal quasi-long-range order which is reflected in the power-law decay of the correlation function
@@
and is plotted in Fig. 4 (blue squares) and this is a signature of the SF phase in one dimension. For comparison, we also show the exponential decay of $\Gamma(r)$ for $V=-2.0$ (green circles) for which the system is gapped.
Figure 5: (a) and (b) display the bond energy $\langle B_{i}\rangle$ and on-site particle density $\langle n_{i}\rangle$ for the points $V=2.0$ and $V=5.6$ marked as red star and blue square in the phase diagram Fig. 2, respectively. The finite oscillations in the bond energy for $V=2.0$ signify the presence of the bond ordering in the system , while the finite on-site density oscillations indicate the presence of a CDW ordering exhibiting a density distribution of $00110011...$ type.

### v2

@@ Abstract
We investigate the phase diagram of a one-dimensional model of hardcore bosons or spinless fermions with tunable nearest-neighbor interactions. By introducing alternating repulsive and attractive interactions on consecutive bonds, we show that the system undergoes a quantum phase transition from a bond-ordered (BO) phase to a charge-density wave-II (CDW-II) phase as the attractive interaction strength increases at a fixed repulsive interaction. For a specific interaction pattern, the BO phase exhibits topological properties, which vanish when the pattern is altered, leading to a transition from a topological BO phase to a trivial BO phase through a gap-closing point where both interactions vanish. We identify these phases using a combination of order parameters, topological invariants, edge-state analysis and Thouless charge pumping. By extending our analysis beyond half-filling, we explore the phase diagram across all densities and identify the superfluid (SF) and the pair-superfluid (PSF) phases, characterized by single-particle and bound-pair excitations at incommensurate densities. The proposed model is experimentally realizable in platforms such as Rydberg excited or ultracold atoms in optical lattices, offering a versatile framework to study such interplay between topology and interactions in low-dimensional systems. Finally, we present the complete ground state phase diagram for all possible signs of the interaction strengths, repulsive or attractive, on consecutive bonds.
@@ I Introduction
The study of topological phases and phase transitions has been a subject of intense research in condensed matter physics Hasan and Kane (2010). Primarily based on a single-particle picture, the topological phases or topological insulators have shaped our understanding of solid-state physics by introducing a new framework for classifying condensed matter systems. The central feature that distinguishes topological phases from non-topological ones is the bulk-edge correspondence, which links the bulk topological invariants to the existence of robust edge or interface states at the boundaries between topologically distinct phases. These edge states are protected by certain underlying symmetries and remain stable against small perturbations until the bulk gap remains finite. Such phases are known as the symmetry-protected topological (SPT) phases Senthil (2015); Rachel (2018). The robustness of the SPT phases makes them highly promising for applications in fields like topological quantum computing, scattering-free wave transport, and several other advanced technologies von Klitzing (2017); Senthil (2015); Fidkowski and Kitaev (2011); Pollmann et al. (2010); Gu and Wen (2009); Chen et al. (2012); Pachos and Simon (2014); Kumar et al. (2023); Zhang and Zhou (2017); Maffei et al. (2018).
@@
One of the celebrated models that possesses such a non-trivial gapped SPT phase is the one-dimensional Su-Schrieffer-Heeger (SSH) model Su et al. (1979) where electrons experience dimerized nearest-neighbor hoppings. The appearance of such unconventional phenomenon in a simple one-dimensional setting has led to numerous exploration of the topological phases in systems described by the SSH model and its variants Asbóth et al. (2016a); Padavić et al. (2018); Anastasiadis et al. (2022); Li et al. (2014); Pérez-González et al. (2019); Ahmadi et al. (2020). Recent progress in the experimental front has propelled the observation of the topological phases in various quantum simulators such as cold atoms in optical lattices, photonic lattices, superconducting circuits, trapped ion arrays, mechanical systems, acoustic systems, electrical circuits, and Rydberg atomic set-ups Atala et al. (2013); Nakajima et al. (2016); Lohse et al. (2016); Mukherjee et al. (2017); Lu et al. (2014); Thatcher et al. (2022); Meier et al. (2016); Xie et al. (2019); Kitagawa et al. (2012); Leder et al. (2016); de Léséleuc et al. (2019); Le et al. (2020); Jörg et al. (2024); Liu et al. (2025); Wu et al. (2021).
@@
Topological phases in non-interacting systems have yielded a wealth of insights and established a robust framework based on band topology and topological invariants. However, interacting systems provide a more challenging domain owing to the competing effects of symmetries, strong correlations, lattice topology, and particle statistics Rachel (2018). This has motivated numerous studies which have explored the effects of inter-particle interactions on the stability of topological phases arising from the SSH-type models Li et al. (2013); Taddia et al. (2017); Grusdt et al. (2013); Manmana et al. (2012); Yoshida et al. (2018); Ryu and Hatsugai (2002); Di Liberto et al. (2016, 2017); Nakagawa et al. (2018); Fraxanet et al. (2022); Wang et al. (2015); Barbiero et al. (2018); Ye et al. (2016); Kuno (2019); Montorsi et al. (2022); Julià-Farré et al. (2024); Ryu and Hatsugai (2002); Delplace et al. (2011); Zhou et al. (2023); Nersesyan (2020); Melo et al. (2023); Chitra et al. (1995); Mohamadi and Abouie (2025). Interestingly, recent studies have shown that strong interactions under proper conditions can favor topological phases  Li and Haldane (2008); Julià-Farré et al. (2022); Argüello-Luengo et al. (2024); Dalla Torre et al. (2006); Rossini and Fazio (2012); Berg et al. (2008); Greschner et al. (2020); Padhan et al. (2024). Even more fascinating is a scenario where interactions alone induce topological phases in systems that are completely trivial in the non-interacting limit Mondal et al. (2022a); Hetényi (2020); Parida et al. (2024); Nigam et al. (2025); Zuo et al. (2020). Recently, several such interaction effects and the interaction-induced SPT phases have been observed using various experimental platforms Sompet et al. (2022); Li et al. (2023); de Léséleuc et al. (2019); Jürgensen et al. (2023, 2021).
@@
Among the existing experimental platforms the systems of ultracold dipolar quantum gases in optical lattices have emerged as versatile platforms for controlling and manipulating the requisite off-site inter-particle interactions which appropriately mimic the models that host topological phases Baier et al. (2016); Sompet et al. (2022). By exploiting the tunability of such interactions, we reveal in this work that in one dimensional lattice, an interaction-induced SPT phase can emerge at half-filling of hardcore bosons or spinless fermions. By assuming uniform hopping amplitudes of the particles (i.e., no dimerization of the hopping) and allowing attractive and repulsive nearest-neighbor (NN) interactions on the alternate bonds of the lattice as depicted in the schematic in Fig. 1(a), we show that the system turns topological for any finite values of the interaction strengths, while the non-interacting limit is a gapless superfluid (SF) or Luttinger liquid (LL). This allows for a phase transition from topological to trivial phase by altering the pattern of the NN interaction (i.e., by changing from attractive-repulsive to repulsive-attractive as depicted in the schematic in Fig. 1(b)) through the point when both the interactions vanish. We demonstrate the signatures of the topological phase and the phase transition through various equilibrium properties. Additionally, we identify further evidence of the topological phase transition via Thouless charge pumping. For completeness, we extend our analysis to explore the situation at all the densities. We also discuss the experimental feasibility of realizing our model. Finally, we present a comprehensive discussion of the ground state phase diagram at half-filling for all possible signs of the interaction strengths.
@@
Figure 1: Schematic illustration of the interaction pattern in NN bonds of the lattice, where ++ denotes repulsive interaction and - denotes attractive interaction.
@@
The structure of this paper is as follows. In Sec. II, we describe the model under consideration and the methodology used to obtain all the results. Sec. III discusses the key findings for the case where the interaction strengths have opposite signs, i.e., repulsive or attractive, on alternate nearest-neighbor bonds. Sec. IV discusses the complete ground state phase diagram for all possible signs of the interaction strengths. Finally, Sec. V summarizes the main results. The details of the analytical calculations are presented in the Appendices.
@@ II Model and Method
We consider a Hamiltonian for a system of interacting hardcore bosons with dimerized NN interactions which is given by
$H$ $=$ $-t~\sum\limits_{i}({a}^{\dagger}_{i}{a}_{i+1}+\text{H.c.})$ (1)
@@
where $i$ denotes the site index, $a_{i}$ ($a^{\dagger}_{i}$) are the annihilation (creation) operators at site $i$, and $n_{i}=a_{i}^{\dagger}a_{i}$ represents the number operator at site $i$. The parameter $t$ corresponds to the hopping amplitude between NN sites, while $V_{1}$ and $V_{2}$ denote the strengths of the dimerized NN interactions. In our numerical calculations presented below, we set $t=1$ which sets the energy scale in the system. The specific form of the NN interaction terms ensures particle-hole symmetry in the Hamiltonian.
@@
where the bosonic occupation state $|0\rangle$ and $|1\rangle$ map to the spin states $|\downarrow\rangle$ and $|\uparrow\rangle$ respectively. The Hamiltonian for the spin chain can then be written as
@@
The Hamiltonian in Eq. (3) can also be mapped to a model of interacting spinless fermions using the Jordan-Wigner transformation Jordan and Wigner (1928) defined as
@@
We have assumed open boundary conditions for this mapping, as will be used throughout this work. With periodic boundary conditions, however, an additional boundary term arises, which depends on the total number of fermions (or total $S_{z}$) in the chain. The above mappings ensure that the models shown in Eqs. (1), (3) and (5) share similar ground state phase diagrams. The results presented below are all obtained by considering the model shown in Eq. (1) which corresponds to a system of hardcore bosons.
The above model has been previously studied using mean-field theory Fisher et al. (1989); Sheshadri et al. (1993); Jaksch et al. (1998); van Oosten et al. (2001), the density-matrix renormalization group (DMRG) White (1992); Schollwöck (2005) method as well as the bosonization approach, for both hardcore bosons and spins  Mondal et al. (2022a); Nigam et al. (2025). These findings highlight that while the non-interacting version of the model remains gapless, introducing a dimerized NN repulsion for hardcore bosons (or dimerized repulsive $ZZ$ exchange interaction for spins) leads to a topological phase and an associated topological phase transition. The topological phase transition occurs as a function of the interaction dimerization which is tuned by fixing the NN interaction strengths for the even (odd) bonds and varying it on the odd (even) bonds of the lattice. This phase transition is marked by a gap-closing point, occurring when the interactions are of equal strengths. However, the scenario considered in this work is completely different as we assume the NN interaction to be attractive in nature in either all the even or odd bonds (see Fig. 1).
@@ III Results
In this section, we discuss our findings in detail. We first present the bulk phase diagram and then analyze the topological properties at half-filling. We also provide the signature of the topological phase through Thouless charge pumping. Subsequently, we analyze the scenario by moving away from half-filling following which we discuss the experimental feasibility of the model. In the end, we briefly discuss the full ground state phase diagram by considering all possible signs of the interaction strengths - both attractive and repulsive.
Figure 2: (a) The phase diagram of the model in the $V_{1}-V_{2}$ plane at half-filling shows the BO0, BOπ and the CDW-II phases. The red circle at the origin depicts the gapless critical point. The diamond symbols indicate the boundary separating the BO and the CDW-II phases. (b) shows a cartoon diagram depicting the interaction pattern considered. (c), (d) and (e) show the schematics for the BO0, CDW-II, and BOπ phases, respectively.
@@ III.1 Phase diagram at half-filling
The ground state phase diagram of the system of hardcore bosons at half-filling is depicted in Fig. 2(a) in the $V_{1}$-$V_{2}$ plane. The upper left (lower right) corner of the figure corresponds to the situation when $V_{1}<0$, $V_{2}>0$ ($V_{1}>0$, $V_{2}<0$). For each case, we numerically find that the entire phase diagram is gapped and there is a gapped-to-gapped transition. In the following, we identify the two gapped phases as the bond order (BO) and the charge-density wave (CDW) phases except at the origin where the system is gapless (denoted by a red circle).
Figure 3: Single-particle excitation gap $G_{1}$ and the bulk gap $G$ are plotted as a function of $V$ along the dashed line drawn in Fig. 2. $G$ (solid curve) remains finite for all the values of $V\neq 0$. $G_{1}$ vanishes in the region $4.3\lesssim V\leq 0$ and remains finite else where. The insets show the finite-size extrapolation of gaps $G_{1}$ (lower panel) and $G$ (upper panel) for three exemplary parameter values corresponding to the gapless ($V=0$) and the gapped phases (V=2.0V=-2.0 and V=6.0V=-6.0). In the inset, the solid lines are the fitted curves and the symbols represent the numerical data.
We find that for any finite value of $V_{2}$ when $V_{1}=0$, the ground state is a BO phase (the upper left part of the phase diagram). However, as $V_{1}$ is increased on the attractive side, a transition to a CDW phase occurs. This results in a gapped-to-gapped transition as a function of attractive $V_{1}$ for each value of repulsive $V_{2}$. A similar situation is found upon interchanging the nature of the interactions i.e., by making $V_{1}$ repulsive and $V_{2}$ attractive; the corresponding phase diagram is shown in the bottom right corner of Fig. 2(a). In the following, we quantify these phases by computing different relevant physical quantities as a function of $V_{1}=-V_{2}=V$ chosen for simplicity. This corresponds to an exemplary cut drawn in the phase diagram which is denoted by a dashed line in Fig. 2(a). Along this line, when $V<0$, all the odd bonds experience attractive interaction while the even bonds experience repulsive interaction due to the boundary condition chosen (i.e. OBC). Conversely, for $V>0$, the odd bonds become repulsive, and even bonds become attractive.
Figure 4: The change in density ($\rho$) as a function of change in chemical potential ($\mu$) for different values of $V$ for a system size $L=80$. The plateaus at half-filling ($\rho=1/2$) indicate the gap in the system and the length of the plateaus give the estimate of the gaps.
Figure 5: The correlation function $\Gamma(r)$ plotted as a function of distance (r=|ij|r=|i-j|) for a system size $L=200$ showing power-law and exponential decays for $V=0.0$ and V=2.0V=-2.0, respectively. The correlation values taken from $L/4$ to $3L/4$ to avoid the edge effects.
Figure 6: (a) and (b) display the bond energy $\langle B_{i}\rangle$ and the on-site particle density $\langle n_{i}\rangle$ as a function of $i$ for $V=1.0$ and $V=5.6$ marked as red star and blue square, respectively in the phase diagram shown in Fig. 2. The finite oscillations in the bond energy for $V=1.0$ signify the presence of the bond ordering in the system, while the finite on-site density oscillations indicate the presence of a CDW ordering exhibiting a density distribution of $\cdots 11001100\cdots$ type.
[…]
@@
$\mu^{+}_{1}=E_{N+1}-E_{N}~~~\text{and}~~~\mu^{-}_{1}=E_{N}-E_{N-1},$ (7)
are the chemical potentials corresponding to adding and removing a particle respectively in the system, and $E_{N}$ represents the ground state energy of the system with $N$ bosons. In Fig. 3, we plot the extrapolated values of the gap $G_{1}$ (dotted plus curve) as a function of $V_{1}=-V_{2}=V$. It can be seen that the gap remains finite in the region $V>0$. However, in the region $V\leq 0$, the gap $G_{1}$ vanishes up to a critical $V\sim-4.3$, after which it becomes finite. In the following, we will show that while the vanishing of $G_{1}$ at $V=0$ is due to the gapless nature of the system at that point, for $4.3\lesssim V<0$, the zero gap is due to the topological nature of the system. In order to show that the system is gapped in all the other parameter regimes except $V=0$, we examine the change in density ($\rho$) as a function of the chemical potential ($\mu$)  Hikihara et al. (2008); Mishra et al. (2015a, b) for different values of $V$ which is shown in Fig. 4. The density is defined as $\rho=N/L$, where $N$ is the total particle number and $L$ is the system size. The gapped nature can be seen as the plateaus in the $\rho-\mu$ curve at half-filling and the length of the plateau represents the gap ($G$) in the system. We obtain the gap as
$G=\mu^{+}-\mu^{-},$ (8)
where $\mu^{+}$ and $\mu^{-}$ are the chemical potentials corresponding to the right and left ends of the plateau for each curve. The extrapolated values of $G$ (red solid curve) as a function of $V$ are plotted in Fig. 3 where the finite values of $G$ for $V\neq 0$ confirm the gapped nature of the phases. At the gapless point $V=0$, the system exhibits off-diagonal quasi-long-range order which is reflected in the power-law decay of the correlation function
@@
and is plotted in Fig. 5 (blue squares). This is a signature of the SF phase in one dimension. For comparison, we also show the exponential decay of $\Gamma(r)$ for V=2.0V=-2.0 (green circles) for which the system is gapped.
We now turn to quantify the gapped phases. As mentioned already, these are the BO and the CDW phases. The BO phase is characterized by finite dimerization of the particles in the NN bonds. To quantify this feature, we use the bond kinetic energy $\langle B_{i}\rangle$, where $B_{i}=a_{i}^{\dagger}a_{i+1}+\text{H.c.}$, exhibits finite oscillations as a function of the bond index $i$, as shown in Fig. 6 (a) for $V=1.0$ (the point marked as a red star in Fig. 2). At the same time, the CDW order manifests itself as finite oscillations in the real-space density $\langle n_{i}\rangle$ which has the form $\cdots 11001100\cdots$ as depicted in Fig. 6 (b) for $V=5.6$ (the point marked as a blue square in Fig. 2). This pattern of the density distribution is the signature of a CDW-II phase.
$\text{S}_{\text{BO}}(k)=\frac{1}{L^{2}}\sum_{i,j}e^{ik|i-j|}\langle B_{i}B_{j}\rangle,$ (10)
and
$\text{S}(k)=\frac{1}{L^{2}}\sum_{i,j}e^{ik|i-j|}\langle n_{i}n_{j}\rangle,$ (11)